OpenStudy (anonymous):

Find x if the average of 20,20,19,13 and x is 16.

OpenStudy (anonymous):

average = summation of numbers / total items so average is given 16 so, 16= (20+20+19+13+x)/ 5 => 16= (72+x) / 5 multiply both sides by 5, 16 x 5=\[\frac{ (72+x)5 }{ 5 }\] 90 = 72+x subtract 72 from both sides, 90-72=72-72+x 18=x so x=18

OpenStudy (anonymous):

thank you!!!